OpenStudy (anonymous):

What is the equation of the line that passes through the points (–12, –8) and (–17, –16)?

OpenStudy (anonymous):

Try using point slope form.

OpenStudy (anonymous):

y−y1=m(x−x1)plug in the values and solve for "y"↑point-slope form

OpenStudy (anonymous):

y-(-16)=8/5x(x-(-17))

OpenStudy (anonymous):

now solve for y

OpenStudy (anonymous):

what kind of equations : vector , paramatic or Cartesian ?

OpenStudy (anonymous):

Cartesian form Y-Y1/X-X1= Slope\[\frac{ Y-Y1 }{ X-X1 }=\frac{ Y2-Y1 }{ X2-X1 }\]